10 years ago, the average age of a family of 4 members was 24 years. Two children having been born (with age difference of 2 years), the present average age of the family is the same. The present age of the youngest child is
Answer by greenestamps(13334)



The average age of the 4 10 years ago was 24; the sum of their ages then was 24*4=96.

Each of them has aged 10 years; now the sum of their ages is 96+4*10 = 136.

The average of all 6 family members now is again 24; so the sum of all 6 ages is 24*6=144.

The difference between 144 and 136 is 8; that represents the sum of the ages of the two youngest children, with one of them 2 years older than the other.

You can finish the problem from there.